mrs. Corazon is 40 years old and his daughter kris is 12. when will the mother be twice as old as the daughter?

Answer by checkley79(3341)

40+X=2(12+X)
40+X=24+2X
40-24=2X-X
16=X ANS. FOR THE ADDITIONAL 16 YEARS BEFORE MRS IS TWICE KRIS'S AGE.
PROOF:
40+16=2(12+16)
56=2*28
56=56